a particle of mass m and carrying charge -q1 is moving around +q2 along a circular path of radius r. Find period of revolution of the charge -q1

Grade:12

2 Answers

Eshan
5 years ago
We are assuming that the charge+q_2is fixed,

The centripetal acceleration of the charge-q_1will be provided by the coulombic attraction by the other charge.
\implies \dfrac{mv^2}{r}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q_1q_2}{r^2}
\implies v=\sqrt{\dfrac{q_1q_2}{4\pi m\epsilon_0r}}
Hence the time period of revolution=\dfrac{2\pi r}{v}=\sqrt{\dfrac{16\pi^3 m\epsilon_0 r^3}{q_1q_2}}
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
The coulombic force will be the source of centripetal force, hence both are equal.
On solving this we get, Period of revolution, T = 16π3mε0r3/q1q2
 
Thanks and regards,
Kushagra